OpenStudy (anonymous):

hi advance question related to definite integral , i know how to do it but just not sure ?

OpenStudy (anonymous):

i do know answer for i and ii

OpenStudy (unklerhaukus):

what were these results you got for (i) and (ii)?

OpenStudy (anonymous):

i is just show that

OpenStudy (anonymous):

and ii , I got -1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i dunno how to expand it properly

OpenStudy (anonymous):

i have done my solution

OpenStudy (anonymous):

but not sure

ganeshie8 (ganeshie8):

Looks good, In the last line, maybe just leave the answer in exact form : \[\dfrac{\lim\limits_{n\to\infty}~~e^{1/n}(1-e)}{\lim\limits_{n\to\infty}~~n(1-e^{1/n)}}=\dfrac{1(1-e)}{-1}=e-1\]

OpenStudy (anonymous):

ok thank u appreciate it @ganeshie8

ganeshie8 (ganeshie8):

np :)